Vickrey Technology has had net income of $1,500,000 in the current fiscal year. There are 1,000,000 shares of common stock outstanding along with convertible bonds, which have a total face value of $8 million. The $8 million is represented by 5,000 different $1,000 bonds. Each $1,000 bond owes and pays 4% interest. The conversion ratio is 30. The firm is in a 21% tax bracket. What is Vickrey's "diluted earnings per share?"
Answer #1

Calculation of diluted earning per share:

Interest expenses = 5000 * $1000 * 4% = $200,000.

Net income before tax = Net income * 100 / (100 - tax rate)
= $1,500,000 * 100 / (100 - 21)
= $1,898,734.18

EBIT = Net income before tax + interest expense
= $1,898,734.18 + $200,000
= $2,098,734.18

If total convertible bonds converted into shares, then
Net income = EBIT * (1 - Tax rate)
= $2,098,734.18 * (1 - 0.21)
= $1,658,000

Number of shares diluted = 1,000,000 + (5000 * 30)
= 1,000,000 + 150,000
= 1,150,000 shares

Diluted earning per share = Net income / Number of shares
= $1,658,000 / 1,150,000
= $1.44

Vickrey's diluted earning per share = $1.44